22:56
Run 20 started - beam on target
Beam energy is 1.5 MeV. Horizontal beam position = 7.0 mm, vertical beam position = 0.0 mm.
Target density ~ 1e11 at/cm2
Beam intensity at inj ~ 4e7
Detectors are positioned IN. Left arm = 62.71, right arm = 64.60 mm
Histograms and stats zero'd, monitoring code restarted.
Beam in the ring ~ 3minutes.
Pressure in YR09 E target ~ 4.5e-11 mbar.
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
01:15
Target density has increased slighty: ~ 4e11 at/cm2 (update: this only lasted about 5 min)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
3:10
Target density has decresed slightly but will come back (hopefully) ~7e10 at/cm2.
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
05:15
Target density has been somewhat unstable, between ~7E10 and ~3E12 at/cm2.
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
07:37
Run continues. Target density still fluctuating around 1E11 atoms/cm2.
Temperatures OK. Merger OK. Rates OK.
System wide checks as before.
Waveforms OK. See attachments.
Some screenshots from the online analysis. We do not see a peak at 1 MeV. We do however clearly see more D than H2+.
Stopping run to make some tests.
|